# RF-3DGS: Wireless Channel Modeling with Radio Radiance Field and 3D Gaussian Splatting

### v1

@@ Abstract
Precisely modeling radio propagation in complex environments has been a significant challenge, especially with the advent of 5G and beyond networks, where managing massive antenna arrays demands more detailed information. Traditional methods, such as empirical models and ray tracing, often fall short, either due to insufficient details or with challenges for real-time applications. Inspired by the newly proposed 3D Gaussian Splatting method in computer vision domain, which outperforms in reconstructing optical radiance fields, we propose RF-3DGS, a novel approach that enables precise site-specific reconstruction of radio radiance fields from sparse samples. RF-3DGS can render spatial spectra at arbitrary positions within 2 ms following a brief 3-minute training period, effectively identifying dominant propagation paths at these locations. Furthermore, RF-3DGS can provide fine-grained Channel State Information (CSI) of these paths, including the angle of departure and delay. Our experiments, calibrated through real-world measurements, demonstrate that RF-3DGS not only significantly improves rendering quality, training speed, and rendering speed compared to state-of-the-art methods but also holds great potential for supporting wireless communication and advanced applications such as Integrated Sensing and Communication (ISAC).
@@ I Introduction
I-A Modeling the Radio Propagation
Wireless communication systems facilitate the exchange of information carried by Electromagnetic (EM) waves between physically separated Transmitter (Tx) and Receiver (Rx). As EM waves propagate from the Tx to the Rx, they undergo various interactions, including reflection, diffraction, refraction, and scattering. Consequently, these waves may reach the Rx via multiple paths, each characterized by distinct CSI. Radio propagation modeling, which involves characterizing these propagation paths, has been the focus of extensive research due to its crucial role in the design, assessment, and deployment of radio networks [1].
Over the years, various methods for radio propagation modeling have been developed. The empirical models [2, 3, 4, 5], although effective for providing coarse-grained path loss information over large distances (from hundreds of meters to tens of kilometers), lack precision. Computational Electromagnetics (CEM) methods [6, 7], while being powerful for small-scale modeling, such as antenna design and near-field communication [8, 9], are impractical for a wider range of applications. As 5G and emerging 6G networks demand higher accuracy and more comprehensive radio propagation modeling, ray tracing offers a potential solution by approximating radio waves using ray concepts [10, 11, 12, 13, 14]. However, ray tracing is still limited by its high computational complexity and stringent environmental data requirements, making it unsuitable for more general and real-time applications.
@@ I-A Modeling the Radio Propagation
For modeling radio wave propagation using ray tracing methods, a similar trend has emerged in the Computer Vision (CV) domain for 3D scene reconstruction, where volume rendering techniques [15] are employed based on photographs to optimize scene representations [16, 17, 18, 19, 20, 21]. These representations typically capture the optical radiance emitted from object surfaces towards the camera, thereby enabling novel view synthesis and environment reconstruction. In this process, the radiance field is modeled as rays passing through each pixel to the camera, similar to how wireless channels are modeled by Multi-Path Components (MPCs) identified by their Angles of Arrival (AoA). This similarity suggests that such CV techniques could offer valuable insights for enhancing radio wave propagation modeling.
A pioneering work of them is Neural Radiance Fields (NeRF) [16], which achieves exceptional synthesis quality with very sparse input. NeRF’s success is grounded in two key components: differentiable volume rendering and neural radiance representation. Volume rendering visualizes volumetric data by casting rays through pixels and integrating their contributions along the ray to determine each pixel’s color. The differentiability of this process enables gradient descent optimization from training images to the neural radiance representation, which consists of a geometry density Multi-Layer Perceptron (MLP) and a radiance feature MLP. Despite its groundbreaking results, NeRF has notable limitations, including long training times (up to tens of hours), slow inference times (several seconds), and its implicit neural network representation, requiring exhaustive querying of the scene to obtain geometric information.
Recently, 3D Gaussian Splatting (3DGS) [22] has emerged as a trendsetter, achieving higher reconstruction quality with significantly reduced training times (tens of minutes) and much higher rendering speeds (exceeding 100 FPS). This method employs 3D Gaussian distributions (hereafter referred to as “Gaussians”) for geometry and spherical harmonics (SH) functions for radiance. Moreover, it introduces a rasterization pipeline that extensively accelerates both the training and rendering processes. Over the past year, substantial research has built upon 3DGS, exploring applications such as city-scale reconstruction [23], Structure-from-Motion (SfM) free 3D Gaussian Splatting [24], and 4D Gaussian Splatting with additional time domain [25].
@@ I-B RF-3DGS
Returning to our context of radio propagation, the rays that render pixels of the image in the optical radiance field are equivalent to the MPCs that reach the Rx. Consequently, the pixel intensity corresponds to the MPC gain, and the image itself is equivalent to the radio AoA-gain spatial spectrum.
Inspired by such equivalence and the 3DGS method, we propose developing a similar approach for modeling radio wave propagation. However, radio waves differ significantly from visible light. Firstly, in wireless communication, we are concerned not only with the AoA-gain spectrum but also with other CSI, such as Angle of Departure (AoD) and delay of each path [26]. Secondly, due to the physical characteristics of radio waves, there is often a lack of detailed object geometric information in the ground truth radiance field, and limited spatial resolutions of radio devices.
@@
In this paper, we propose RF-3DGS, a method that addresses the aforementioned challenges and enables rapid, precise, site-specific reconstruction of the Radio Frequency (RF) radiance field from very sparse training inputs. In addition to the inherited advantages of 3DGS, such as ultra-fast training and inference speeds, RF-3DGS distinguishes itself from visual 3DGS by not only reconstructing the last-bounce-only radiance field (AoA), but also encoding other CSI, such as Angle of Departure (AoD) and delay, into the SH functions, effectively addressing the first challenge. This enhanced radiance field representation allows for the efficient implementation of angle-domain management in massive Multiple Input Multiple Output (MIMO) systems, thereby avoiding the increasing computational complexity associated with pilot-based estimation as array size grows [27, 28]. Furthermore, this representation’s real-time inference and explicit geometry make it beneficial for other wireless system tasks, such as sensing and digital twin applications [29, 30, 31].
To address the second challenge, we explore various array signal processing techniques and propose a two-stage fusion training strategy that enables more precise geometry representation. Additionally, our method is specifically designed for massive MIMO systems, which demand the high-accuracy spatial spectra that RF-3DGS provides. These systems typically operate in the upper mid-band (7 to 16 GHz) or mmWave bands, where higher frequencies capture more geometric detail and enable larger array sizes, thereby offering significantly higher spatial resolution.
@@
To demonstrate the performance of RF-3DGS, several test examples are presented in Fig. 1. Whether operating with the practical Minimum Variance Distortionless Response (MVDR) spectra generated by an $8\times 8$ patch antenna array or with ideal radio spatial spectra, which require more advanced and costly systems, RF-3DGS effectively reconstructs the radio radiance field in unvisited locations with very sparse input (only tens of samples in a complex environment). Moreover, RF-3DGS exhibits strong extrapolation capabilities thanks to the two-stage fusion training, which integrates visual geometric information into the representation. The effectiveness of RF-3DGS is evident in Fig. 1, where our method significantly outperforms the State-of-The-Art (SOTA) NeRF2 [32].
@@
Quantitatively, RF-3DGS achieves a substantially better Learned Perceptual Image Patch Similarity (LPIPS) [33] value in the reconstructed spatial spectra compared to NeRF2, with an error reduction of 84.64%. LPIPS evaluates similarity between images by comparing deep features extracted from pretrained neural networks, providing a more robust measure of visual quality than other statistical metrics. Meanwhile, RF-3DGS requires only 3 minutes for RF radiance field reconstruction and just 2 ms to render a spectrum at an arbitrary array position, whereas NeRF2 takes 3 hours of training and approximately 1 second to render a new spectrum.
@@
To further evaluate its effectiveness in wireless communication, the synthesized spatial spectra are utilized to generate Conventional Beamforming (CBF) steering vectors. The results indicate that in the ideal spectrum scenario, RF-3DGS accurately guides the steering angle with a median deviation of 5.94 at the test points. Moreover, since RF-3DGS also encodes additional CSI features, the AoD and delay channel spectra achieve LPIPS values of 0.395 and 0.382, respectively. We also assessed our digital twin framework by comparing it with physical twin measurements, which demonstrate a high degree of consistency with the physical reality.
@@ I-C Contributions
  • We utilize 3D reconstruction method from CV to wireless channel modeling, which can “render” 3D CSI at arbitrary locations in real-time (2 ms).
  • We analyze differences between visible light and radio waves affect radiance field reconstruction, propose strategies to mitigate these effects, and explore various array signal processing techniques.
@@
  • We propose the RF-3DGS framework, which addresses two key challenges: 1) representing comprehensive CSI within a radiance field using CSI-encoded SH functions, and 2) fusing visual and radio data into the radiance field representation through a novel two-stage fusion training strategy. The whole pipeline only takes 3-minute in training.
  • We construct a digital twin with a comprehensive radio spatial spectrum simulation pipeline and validate it using real-world measurements. Additionally, we demonstrate an ISAC scenario within this digital twin framework.
II Primer on Radio Radiance Field
In this section, we explore the rationale behind applying optical radiance field reconstruction methods to the RF domain. We will also highlight the critical factors that must be considered to ensure the successful adaptation of these methods.
Figure 2: RF and visible light wave comparison.
II-A Visible Light and Radio Frequency: A Comparison
Although RF waves and visible light are both EM waves, their physical characteristics differ significantly. Firstly, in complex environments, the propagation behaviors of these two types of EM waves are markedly different, as illustrated in Fig. 2. For visible light, due to its short wavelength, most everyday surfaces do not meet the Rayleigh criterion for smoothness, which requires that the surface roughness height $h$ be smaller than $\lambda/(8\cos\theta)$ (where $\lambda$ is the wavelength and $\theta$ is the incidence angle). Consequently, diffusion and scattering become the dominant interactions between visible light and surrounding objects, making these objects visible from almost all directions.
In contrast, for RF waves, especially those in the sub-6GHz range, specular reflection and diffraction are dominant [34, 35]. These interactions exhibit much higher directivity compared to scattering. While such interactions are favorable for wireless communication, providing signal coverage in Non-Line-of-Sight (NLoS) scenarios, they are less effective for ”illuminating” objects. This means that the received radio waves contain less geometric information, which is crucial for the effectiveness of radiance field reconstruction methods.
Fortunately, as wireless communication frequencies extend into the mmWave range, surfaces tend to appear rougher, resulting in more scattering [36]. This enhances the opportunities to extract geometric information from received mmWave signals. Applications of mmWave imaging in surveillance and defense applications further demonstrate this characteristic [37, 38, 39, 40].
Secondly, the short wavelength of visible light allows the use of lenses to perfectly separate incoming rays based on their AoA to the millions of photodiodes in a handheld camera. For RF, constructing such an “optical” imager is physically challenging due to the limited number of effective elements within a constrained aperture. However, antenna arrays offer the advantage of capturing the response of each antenna element, including phase information, thereby enabling the use of array signal processing techniques to separate incoming RF signals and generate the radio spatial spectrum. This spectrum, representing signal strength at different AoAs, can be considered the “RF picture” captured by the antenna array. The quality of this picture, however, is still limited by the physical aperture size, the processing techniques, and the measurement devices.
II-B Radiance Field Reconstruction Methods
Radiance Field Reconstruction methods, including 3DGS, are based on volume rendering, which is typically used for rendering transparent volumes such as fog and fire but can also be adapted to render solid objects due to its flexibility. Fig. 3 illustrates the basic concept of volume rendering. The rendering equation can be expressed as:

### v2

@@ Abstract
Precisely modeling radio propagation in complex environments has been a significant challenge, especially with the advent of 5G and beyond networks, where managing massive antenna arrays demands more detailed information. Traditional methods, such as empirical models and ray tracing, often fall short, either due to insufficient details or because of challenges for real-time applications. Inspired by the newly proposed 3D Gaussian Splatting method in the computer vision domain, which outperforms other methods in reconstructing optical radiance fields, we propose RF-3DGS, a novel approach that enables precise site-specific reconstruction of radio radiance fields from sparse samples. RF-3DGS can render radio spatial spectra at arbitrary positions within 2 ms following a brief 3-minute training period, effectively identifying dominant propagation paths. Furthermore, RF-3DGS can provide fine-grained Spatial Channel State Information (Spatial-CSI) of these paths, including the channel gain, the delay, the angle of arrival (AoA), and the angle of departure (AoD). Our experiments, calibrated through real-world measurements, demonstrate that RF-3DGS not only significantly improves reconstruction quality, training efficiency, and rendering speed compared to state-of-the-art methods, but also holds great potential for supporting wireless communication and advanced applications such as Integrated Sensing and Communication (ISAC). Code and dataset will be available at https://github.com/SunLab-UGA/RF-3DGS.
@@ I Introduction
I-A Wireless Channel Modeling
Wireless communication systems facilitate the exchange of information carried by Electromagnetic (EM) waves between a physically separated Transmitter (Tx) and Receiver (Rx). As EM waves propagate from the Tx to the Rx, they undergo various effects, including reflection, diffraction, refraction, and scattering. Consequently, these waves may reach the Rx via multiple paths, and each path is characterized by its own set of channel properties, known as a multipath component (MPC). In the context of 5G and the upcoming 6G Multiple-Input and Multiple-Output (MIMO) networks, the antenna array size is expected to grow to thousands of elements. This expansion requires the efficient spatial management of such large arrays, thereby mitigating the increasing computational complexity associated with pilot-based estimation as the array size increases [1, 2]. This management, in turn, requires the Spatial Channel State Information (Spatial-CSI) [3], which include not only the traditional channel gain and delay of each MPC, but also the angle of departure (AoD), angle of arrival (AoA), and polarization. Meeting this requirement poses a significant challenge for wireless channel modeling.
Over the years, various methods for wireless channel modeling have been developed. The empirical models [4, 5], although effective for providing coarse-grained path loss information over large distances (from hundreds of meters to tens of kilometers), fail to capture the finer details such as the spatial information of the MPCs. Computational Electromagnetics (CEM) methods [6, 7], while being powerful for small-scale modeling, such as antenna design and near-field communication [8, 9], are computationally impractical for a wider range of applications. Ray tracing, positioned between these two approaches, offers a potential solution by approximating radio waves using ray concepts [10, 11, 12]. However, ray tracing is still limited by its high computational complexity and stringent environmental data requirements (e.g., 3D modeling and surface properties), making it unsuitable for more general and real-time applications.
@@ I-A Wireless Channel Modeling
For modeling wireless channel using ray tracing methods, a similar trend has emerged in the Computer Vision (CV) domain for 3D scene reconstruction, where volume rendering techniques [13] are employed based on photographs to optimize scene representations [14, 15, 16]. These representations typically capture the optical radiance emitted from object surfaces towards the camera, thereby enabling novel view synthesis and environment reconstruction. In this process, the optical radiance is modeled as a group of rays passing through each pixel to the camera, similar to how wireless channels are modeled by a group of MPCs arriving at the Rx from different AoAs. This similarity suggests that CV techniques could offer valuable insights for enhancing radio wave propagation modeling.
A recent pioneering work on CV reconstruction is Neural Radiance Field (NeRF) [14], which achieves exceptional synthesis quality with sparse input. Despite its groundbreaking results, NeRF has notable limitations, including long training times (up to tens of hours), slow inference times (several seconds). More importantly, its implicit neural network representation requires exhaustive querying of the scene to obtain geometric information. Recently, 3D Gaussian Splatting (3DGS) [17] has emerged as an alternative, achieving higher reconstruction quality with significantly reduced training times (tens of minutes) and much higher inference speeds (less than 10 ms). Over the past year, substantial research has built upon 3DGS, exploring applications such as city-scale reconstruction [18], Structure-from-Motion (SfM) free 3D Gaussian Splatting [19], and 4D Gaussian Splatting, which includes the time domain [20].
@@ I-B RF-3DGS
Inspired by the similarity of our problem and scene reconstruction using optical radiance field, we propose developing an approach to model the wireless channel through the Radio Radiance Field (RRF). However, we face the following three challenges: First, since the wireless channel fundamentally differs from photographs, how can we design a learnable RRF representation structure that effectively captures the multi-modal Spatial-CSI? Second, given a well-trained RRF representation, how can we efficiently query the received wireless channel at an arbitrary Rx location? Third, with a set of wireless samples, how can we use them to train the RRF representation to accurately model the wireless channel?
@@
In this paper, we propose RF-3DGS, a method that addresses the aforementioned challenges and enables rapid, precise, site-specific reconstruction of the RRF from very sparse training inputs. Compared to previous works, RF-3DGS offers several advantages, including ultra-fast training and inference speeds, concise representation size, and the ability to reconstruct the Spatial-CSI of each MPC in the RRF reconstruction. This enhanced RRF representation enables efficient spatial management of massive MIMO systems. Furthermore, the real-time inference and explicit geometric representation make it highly beneficial for other wireless system tasks, such as sensing and digital twin applications [21, 22, 23]. In addition, we explore various array signal processing techniques and propose a two-stage fusion training strategy to extract geometric information from visual input and integrate it into the RRF representation.
@@
To demonstrate the performance of RF-3DGS, several test examples are presented in Fig. 1. Whether operating with the practical Minimum Variance Distortionless Response (MVDR) spectra generated by an $8\times 8$ patch antenna array or with ideal radio spatial spectra, which require more advanced and costly systems, RF-3DGS effectively reconstructs the radio radiance field in unvisited locations with very sparse input (only tens of samples in a complex environment). Moreover, RF-3DGS exhibits strong extrapolation capabilities thanks to the two-stage fusion training. The effectiveness of RF-3DGS is evident in Fig. 1, where our method significantly outperforms the State-of-The-Art (SOTA) NeRF2 [24].
@@
Quantitatively, RF-3DGS achieves an 84.64% error reduction in reconstructed spatial spectra quality compared to NeRF2. Meanwhile, RF-3DGS requires only 3 minutes for RF radiance field reconstruction and just 2 ms to render a spectrum at an arbitrary array position, whereas NeRF2 takes 3 hours of training and approximately 1 second to render a new spectrum.
@@
To further evaluate its effectiveness in wireless communication, the synthesized spatial spectra are utilized to generate Conventional Beamforming (CBF) steering vectors. The results indicate that in the ideal spectrum scenario, RF-3DGS accurately guides the steering angle with a median deviation of 5.94 at the test points. Moreover, since RF-3DGS also encodes additional CSI features, the reconstructed AoD and delay channel spectra exhibit high similarity to the ground truth. We also assessed our digital twin framework by comparing it with physical twin measurements, which demonstrate a high degree of consistency with the physical reality.
@@ I-C Contributions
  • First, we propose the fundamentals of the RRF method and theoretically analyze how the RRF can represent radio propagation, how to query the wireless channel from an RRF representation, and how to reconstruct the RRF from wireless samples.
@@
  • Next, we propose the RF-3DGS framework, which is characterized by two key features: (1) representing multi-modal Spatial-CSI using carefully designed CSI-encoded spherical harmonic (SH) functions, and (2) integrating visual and radio data into the RRF representation through a novel two-stage fusion training strategy to address the limitations of wireless samples.
  • Lastly, we construct a digital twin framework comprising comprehensive real-world wireless measurements and precise radio spatial spectra simulations in its digital replica. We further demonstrate how our method benefits massive MIMO spatial management and advanced applications such as Integrated Sensing and Communications (ISAC).
II Fundamentals of the RRF Method
In this section, before detailing RF-3DGS, we first establish the theoretical foundations of the RRF method. We begin by defining the RRF representation structure, which models the radio propagation, followed by a discussion on how the wireless channel can be rendered from this representation. Lastly, we discuss training the RRF representation to accurately model the wireless channel.
II-A Overview of RRF Method
[…]
As shown in Fig. 2, the RRF function $\mathbf{c}(\mathbf{x},\mathbf{d})$ describes the radio radiance field across the entire 3D space. For a specific point $\mathbf{x_{r}}$, the radio radiance $\mathbf{c}_{\mathbf{x_{r}}}(\mathbf{d})$ is a spherical function that describes the MPCs originating or retransmitting from this point at arbitrary direction $\mathbf{d}$. For a specific point $\mathbf{x_{r}}$ and direction $\mathbf{d_{r}}$, they define the geometry of a ray in 3D space, and $\mathbf{c}_{\mathbf{x_{r}},\mathbf{d_{r}}}$ represents the Spatial-CSI values of the MPC corresponding to this ray. Because this ray is actually the final segment of a complete radio propagation path from the Tx to a Rx on this ray, the path also corresponds to an MPC of the wireless channel to the Rx. Thus, hereafter, we sometimes integrate the concepts of a ray and its corresponding MPC as ray (MPC) for simplicity.
For an Rx that lies on a ray (MPC), if the ray is within the Rx’s Field of View (FoV) with no obstacles between the ray origin $\mathbf{x_{r}}$ and the Rx, the Rx can receive this MPC. To handle the potential blockages, the geometry is required. This is represented by the density field function, $\alpha(\mathbf{x})$, which assigns zero to free space, low values to translucent objects, and high values to solid object surfaces (as reconstructing the interior of solid objects is both challenging and unnecessary).
Given a well-trained RRF representation $\mathcal{R}$ and the current Rx configuration, the wireless channel between the Tx and Rx can be rendered as a radio spatial spectrum $\mathcal{S}(\mathcal{R},\mathcal{P_{\mathbf{Rx}}},\mathbf{x_{Rx}},\mathbf{d_{Rx}})$, where $\mathbf{x_{Rx}}$ and $\mathbf{d_{Rx}}$ are the position and direction of the Rx, and $\mathcal{P_{\mathbf{Rx}}}$ is the Rx projection model defined by the Rx array’s FoV and the query requirements. This radio spatial spectrum enables efficient querying of Spatial-CSI for MPCs and facilitates high-level applications.
II-B Structure of the RRF Representation
For the RRF representation =$\mathcal{R}=\{\mathbf{c}(\mathbf{x},\mathbf{d}),\alpha(\mathbf{x})\}$, although the RRF function $\mathbf{c}(\mathbf{x},\mathbf{d})$ can describe the Spatial-CSI of the radio radiance originating from every 3D point $\mathbf{x}$, as shown in Fig. 2, the geometry is still necessary for two reasons. First, as mentioned, the subsequent rendering process requires the geometry to handle the blockages. Second, for efficiency, we only need to reconstruct and render the RRF function $\mathbf{c}(\mathbf{x},\mathbf{d})$ for object points $\mathbf{x_{obj}}$, where $\mathbf{x_{obj}}\in\{\mathbf{x}|\alpha(\mathbf{x})\neq 0\}$. These considerations arise because physical radio radiance is only actually emitted from radiance sources, such as the Tx and retransmitting surfaces depicted by the red dots in Fig. 2. In free space, the RRF function $\mathbf{c}(\mathbf{x},\mathbf{d})$ is assigned with zero value, which should be distinguished from the EM field that spreads the whole space. The latter requires significantly more resources for simulating, reconstructing, and representing.
Figure 2: Represent multi-path propagation of radio waves using the Radio Radiance Field.
More specifically, for a given ray (MPC) defined by its ray origin $\mathbf{x_{r}}$ and direction $\mathbf{d_{r}}$, the Spatial-CSI values $\mathbf{c}_{\mathbf{x_{r}},\mathbf{d_{r}}}$ of this ray (MPC) should include channel gain, delay, AoA, AoD, and polarization. However, in this work, we only encode the channel gain, delay, and AoD into $\mathbf{c}_{\mathbf{x_{r}},\mathbf{d_{r}}}$, as the AoA is already determined by the geometry of this ray and the Rx direction, and the polarization will be considered in future work due to its challenges in field measurement and simulation. Therefore, a practical representation of the RRF is =$\mathcal{R}=\{\mathbf{c}(\mathbf{x_{obj}},\mathbf{d}),\alpha(\mathbf{x})\}$, which encompasses the geometry of the entire 3D space and the radio radiance at each object point.
II-C Rendering the RRF Representation
As mentioned earlier, the purpose of RRF method is rapid and accurate querying of the Spatial-CSI of wireless channel between a fixed Tx and an arbitrary Rx position.
To achieve this, given a well-trained RRF representation $\mathcal{R}$, the goal is to query how radiance arrives at a Rx position from different directions. As shown in the upper-left part of Fig. 3, a practical approach is to query the Spatial-CSI at a set of uniformly distributed AoAs (gray arrows) within the array’s FoV (red spherical cap), assuming the presence of MPCs at each AoA, even though most queried AoAs may not exhibit strong MPCs. The queried Spatial-CSI values then form the radio spatial spectrum $\mathcal{S}(\mathcal{R},\mathcal{P_{\mathbf{Rx}}},\mathbf{x_{Rx}},\mathbf{d_{Rx}})$, which accurately depicts the wireless channel between the Tx and Rx.
Figure 3: Render a radio spatial spectrum from the RRF representation.
In particular, the projection model $\mathcal{P_{\mathbf{Rx}}}$ defines how the AoAs to be queried are distributed within the FoV and how the Spatial-CSI values are mapped to pixels in a 2D image, similar to an “RF picture”. These “RF pictures” have parallel channels corresponding to mentioned channel gain, delay, and AoD, much like RGB images, but require more careful processing. The design of the projection model $\mathcal{P_{\mathbf{Rx}}}$ is flexible. For better visualization, we use a simple pinhole camera model, consisting of a 200 $\times$ 300 pixel grid with a horizontal FoV of 90, as shown in Fig. 1. More efficient projection model will need to be investigated for future practical implementations.
For querying at each AoA, the alpha-blending rendering technique is applied. As shown in Fig. 3, given a AoA to be queried, we launch a rendering ray along this AoA. This rendering ray is represented as $\mathbf{r}(t)=\mathbf{x_{Rx}}+t\mathbf{d_{ray}}$, where the 3D position $\mathbf{x}$ is simplified to a one-dimensional coordinate $t$ on the ray for convenience. $\mathbf{d_{ray}}$ is the direction of this rendering ray from the array central point $\mathbf{x_{Rx}}$ to the red surface point. To render the received Spatial-CSI $\mathbf{c_{recv}}$ (should be distinguished from the RRF function), the alpha-blending rendering equation can be expressed as: